OpenStudy (anonymous):

A club decides to sell T-shirts for $12 as a fund-raiser. It costs $20 plus $8 per T-shirt to make the T-shirts. Write and solve an equation to find how many T-shirts the club needs to sell in order to make a profit of at least $100.

OpenStudy (anonymous):

12X - (20 + 8X) = 100 solve for X

OpenStudy (anonymous):

Do you by any chance know the equation to how many T-shirts the club needa??

OpenStudy (anonymous):

I have written equation in terms of X see.. simplifying it becomes 4X-20= 100... 4X=120 X=30

OpenStudy (anonymous):

it says at least $100....so this is supposed to be an inequality. swarup did it correctly except it should be : 12x - (20 + 8x) >= 100 12x - 20 - 8x >= 100 4x >= 120 x >= 30 The club would have to sell 30 or more shirts to make a profit of $100 all is the same as swarup, except be sure and put an inequality sign instead of just an equal sign.